I receive a $1,000,000 inheritance on January 1st, 2020. I cannot withdraw any of it until January 1st, 2035. I plan on withdrawing the same amount for 30 years. How much can I withdraw each year? (assuming an 8% interest rate)

****Please explain how you got the answer. I need to learn how to do this****
Answer #1

ANSWER

Annual withdrawal can be $ 281,775.64

Step-1:Future value of $ 1,000,000 on January 1st 2035
Future value = P*(1+i)^n Where,
= 1000000*(1+0.08)^15 P = $       10,00,000
= $ 31,72,169.11 i = 8%
n = 15
Step-2:Calculation of annual withdrawal for 30 years
Annual withdrawal = Investment amount at the beginning of January 1st 2035 / Present value of annuity of 1
= $ 31,72,169.11 / 11.25778
= $ 2,81,775.64
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.08)^-30)/0.08 i = 8%
= 11.25778334 n = 30

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************